# The Multi-Query Paradox in Zeroth-Order Optimization

### v1

@@ 4.2 Optimal Query Allocation
Let $f$ be an $L$-smooth and convex function. The ZO-Avg algorithm with step size $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$ achieves the following convergence guarantee after $T$ iterations:
@@
Under a fixed total query budget $K$, this convergence rate is optimized by choosing $q_{t}=1$ for all t=0,,T1t=0,...,T-1, resulting in $T=K$ iterations and a rate of:
@@
See Appendix A.1. ∎
Theorem 4.3 reinforces the conclusion from the strongly convex case: for the ZO-Avg estimator, using multiple queries per step is an inefficient use of the query budget. Next, we show that for the ZO-Align estimator, the conclusion remains the opposite.
@@
See Appendix A.1. ∎
Comparing the rates for ZO-Avg and ZO-Align, we see a factor of $d+2$ versus $d$. More importantly, the ZO-Align rate depends only on the total number of queries $K$, not on how they are allocated across iterations. This key distinction stems directly from the mathematical properties of the estimators. The per-iteration progress for ZO-Align is linear in the number of queries. However for ZO-Avg, it is a nonlinear concave function.
Non-Convex Case
Then we proceed to non-convex setting, where the objective is only assumed to be $L$-smooth. In this scenario, the goal is typically to find a first-order stationary point (Nocedal and Wright 2006), i.e., a point $x$ where $\|\nabla f(x)\|^{2}\leq\epsilon$. The convergence is measured by the number of queries required to achieve this goal. The analysis reveals the same fundamental dichotomy between the two estimators.
@@
See Appendix A.1. ∎
The conclusions for the non-convex setting mirror those from the convex case, reinforcing the same central dichotomy. While the same fundamental difference in optimal allocation holds, the performance gap between the two methods is less pronounced here than in the strongly convex setting, boiling down to a constant factor difference rather than an exponential one.
@@
We now extend our analysis to the stochastic setting, a standard framework in modern machine learning where the objective is an expectation: $f(x)=\mathbb{E}_{\xi}[F(x,\xi)]$. At each iteration $t$, the true function $f(x_{t})$ is inaccessible. Instead, our zeroth-order estimators must be constructed from a single stochastic realization $F(x_{t},\xi_{t})$ using a random sample $\xi_{t}$. Our zeroth-order estimators are then constructed using this stochastic function. We show that the fundamental dichotomy between ZO-Avg and ZO-Align persists even in the presence of this new source of noise.
@@ 4.3 The Multi-Query Paradox
Our theoretical analysis across strongly convex, convex, non-convex, and stochastic functions reveals a consistent and powerful dichotomy. The effective use of multiple queries per step is critically dependent on the method used to aggregate the information from those queries. This suggests that the common pursuit of an optimal intermediate block size $q$ where ($1<q<d$) may be misguided. Instead, the landscape of multi-query zeroth-order optimization is defined by two fundamentally different, opposing poles of operation.
A deeper insight emerges when comparing the performance gap between the two estimators across the different function classes. The superiority of ZO-Align is most pronounced in the strongly convex setting and diminishes as the problem structure weakens to convex and non-convex. This is not a coincidence, but rather a reflection of how the problem’s geometry dictates the value of a high-quality gradient estimate. For strongly convex functions, the analysis relies on a multiplicative decrease in error at each step, where the contraction factor is directly improved by a better gradient estimate. The benefit of ZO-Align’s superior update direction compounds exponentially, leading to a vastly different convergence rate.
In contrast, the analyses for convex and non-convex functions rely on weaker potential-based arguments that result in a telescoping sum over the iterations. This framework captures cumulative, additive progress rather than a compounding, multiplicative effect. Within this additive framework, the step-by-step superiority of ZO-Align is averaged out over the entire optimization run. Its advantage manifests only as a slightly better constant factor in the final rate—a factor that becomes marginal for large $d$. The lack of strong geometric structure prevents the algorithm from fully capitalizing on the superior search direction in a compounding manner.
Our analysis of the stochastic convex case further reinforces the central thesis established in the deterministic setting: leveraging parallel queries in ZO methods collapses to a choice between two canonical algorithms. The introduction of noise from the stochastic oracle does not create a new middle ground; it simply reinforces the existing poles.
This trade-off between statistical efficiency and the cost of queries pushes the optimal strategy for each estimator to an extreme. For the ZO-Avg estimator, our theory proves that the optimal strategy is always to set $q_{t}=1$, which reduces the algorithm to a simple, sequential single-query approach. Conversely, for the statistically superior ZO-Align estimator, the optimal strategy is to use the maximum number of queries, setting $q_{t}=d$. In this important special case, the projection matrix $P_{u}=U(U^{T}U)^{-1}U^{T}$ becomes the identity matrix $I$ for any full rank $U$. The ZO-Align estimator thereby converges to the true gradient, and the method transforms into a classical full-subspace finite-difference method.
Therefore, our analysis suggests that the question of how to leverage multiple queries in ZO methods does not lead to a novel intermediate-query algorithm. Rather, it collapses the problem to a choice between two well-established canonical algorithms: a sequential single-query method (the optimal realization of ZO-Avg) and a full-subspace finite-difference method (the optimal realization of ZO-Align). Our work clarifies that these two poles are the theoretically-grounded optimal choices, and the path taken depends entirely on the chosen aggregation strategy.
5 Experiments
@@
To complement our theoretical analysis, we conduct a series of experiments designed to investigate the practical performance implications of the ZO-Avg and ZO-Align estimators across several classic optimization problems. The results provide a clear, practical demonstration of the query allocation dichotomy that forms the central thesis of our work.
5.1 Experimental Setup
Objective Functions. We evaluate the algorithms on four classic objective functions representing strongly convex, convex, non-convex, and stochastic convex problems, with the dimension fixed at $d=1000$.
  • Strongly Convex: A quadratic function, $f(x)=\frac{1}{2}x^{T}Ax+b^{T}x$.
  • Convex: A logistic regression function on synthetic data, $f(x)=\frac{1}{m}\sum_{i=1}^{m}\log(1+\exp(-y_{i}a_{i}^{T}x))$.
  • Non-Convex: The Rosenbrock function (Rosenbrock 1960), $f(x)=\sum_{i=1}^{d-1}[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}]$.
  • Stochastic Convex: A stochastic logistic regression objective, $f(x)=\mathbb{E}_{(\mathbf{w},y)\sim\mathcal{D}}[\log(1+\exp(-y\mathbf{w}^{T}x))]+\frac{\rho}{2}\|x\|^{2}$. The function is computed using mini-batches of data sampled at each iteration.
Evaluation Metric. We evaluate ZO-Avg and ZO-Align with various query sizes per iteration (q), plotting the objective value against the number of queries used. Each curve represents the average of 10 independent runs for reliability. We test under two total query budgets: a large budget of $K=20,000$ and a constrained budget of $K=500$. The latter case, where the total budget is less than the problem dimension ($K<d$), is crucial for examining the performance dichotomy in a resource-constrained setting where full-subspace estimation is impossible. Other detailed settings are provided in Appendix A.2.
5.2 Empirical Results and Discussion
In this section, we present the empirical results of our experiments, which validate the theoretical dichotomy established in Section 4.3. The following figures illustrate the performance of the ZO-Avg and ZO-Align estimators on various objective functions. For these figures, the horizontal axis represents the number of queries used, while the vertical axis shows the objective value achieved. With the same number of query used, increasing the number of queries per iteration (q) inversely reduces the total number of iterations. To highlight the focus of our research, we use the total query number used rather than the number of iterations to measure the optimization cost, as this more directly reflects the computational efficiency of zeroth order methods.
As shown in Figure 2 for the strongly convex case, ZO-Align with a large query block size is dramatically superior, while ZO-Avg’s performance degrades as $q$ increases. This holds for both the large ($K=20,000$) and small ($K=500$) budgets. The small-budget scenario is particularly revealing: even when the total budget is too small to afford a single full-subspace estimate ($K<d$), ZO-Align with the largest feasible $q$ still outperforms other strategies, highlighting the robust benefit of its aggregation method. For the convex and non-convex problems, Figures 2 and 4 show that this same fundamental pattern persists. However, the performance landscape is more nuanced, as the improvement of the best ZO-Align strategy over the baselines is more modest than in the strongly convex case. This aligns with our theory, which posits that without the compounding effect of strong convexity, the benefit manifests as an improved constant factor rather than an exponential separation.
The stochastic convex setting, shown in Figure 4, further confirms the robustness of our central thesis, as the core dichotomy between the estimators persists even in the presence of noise from the objective function itself. ZO-Align with large $q$ greatly outperform other competitors.
In summary, the experiments confirm that the multi-query optimization challenge collapses to a choice between two fundamentally different algorithms, with the aggregation method as the critical factor. ZO-Avg is an inefficient use of a query budget, making a sequential single-query method its optimal implementation. In contrast, ZO-Align effectively leverages multiple queries, with its performance systematically improving as the query block size increases.

### v3

@@ 4.2.2 Convex Case
Let $f$ be an $L$-smooth and convex function. The ZO-Avg algorithm with step size stated in (3) achieves the following convergence guarantee after $T$ iterations:
@@
Under a fixed total query budget $K$, this convergence rate is optimized by choosing $q_{t}=1$ for all $t$, resulting in $T=K$ iterations and a rate of:
@@
See Appendix B.7. ∎
Theorem 4.4 reinforces the conclusion from the strongly convex case: for the ZO-Avg estimator, using multiple queries per step is an inefficient use of the query budget. Next, we show that for the ZO-Align estimator, the conclusion remains the different.
@@
See Appendix B.8. ∎
The ZO-Align rate depends only on the total number of queries $K$, not on how they are allocated across iterations. This key distinction stems directly from the mathematical properties of the estimators. The per-iteration progress for ZO-Align is linear in the number of queries. However for ZO-Avg, it is a nonlinear concave function.
4.2.3 Non-Convex Case
Then we proceed to non-convex setting, where the objective is only assumed to be $L$-smooth. In this scenario, the goal is typically to find a first-order stationary point (nocedal2006numerical), i.e., a point $x$ where $\|\nabla f(x)\|^{2}\leq\epsilon$. The convergence is measured by the number of queries required to achieve this goal. The analysis reveals the same fundamental dichotomy between the two estimators.
@@ 4.2.3 Non-Convex Case
See Appendix B.10. ∎
The conclusions for the non-convex setting mirror those from the convex case, reinforcing the conclusion of multi-query paradox. While the same fundamental difference in optimal allocation holds, the performance gap between the two methods is less pronounced here than in the strongly convex setting, boiling down to a constant factor difference rather than an exponential one.
@@
We now extend our analysis to the stochastic setting, a standard framework in modern machine learning where the objective is an expectation: $f(x)=\mathbb{E}_{\xi}[F(x,\xi)]$. We show that the fundamental dichotomy between ZO-Avg and ZO-Align persists.
@@ 4.3 The Multi-Query Paradox
Taken together, the theoretical results reveal a simple dichotomy under a fixed total query budget. For ZO-Avg, all four regimes we study—strongly convex, convex, non-convex, and stochastic convex—lead to the same conclusion: the optimal allocation is to take a single query per iteration, so that the budget is spent on as many steps as possible. For ZO-Align, the picture is the opposite in the strongly convex and stochastic convex cases, where the best strategy is to use as many queries as possible per step (up to $d$ when feasible), while in the deterministic convex and non-convex cases the convergence rate depends only on the total number of queries and is essentially indifferent to how they are distributed across iterations.
This dichotomy has a clear geometric interpretation. ZO-Avg forms its update by averaging directional derivatives along a block of sampled directions, but it never accounts for the geometry of these directions. If several query vectors happen to be close to each other, their information is effectively counted multiple times, while directions that would explore new parts of the space may be underrated. Additional queries therefore provide diminishing returns: they mainly reduce variance through repeated measurements along nearly redundant directions, and this variance reduction scales sublinearly in the block size. Once the linear cost in queries is taken into account, it is always better for ZO-Avg to take more, cheaper steps with a single direction per step than to aggregate many directions at once.
By contrast, ZO-Align uses the entire block of directional measurements to reconstruct the orthogonal projection of the gradient onto the span of the sampled directions. The Gram matrix is explicitly inverted, so correlations and unequal norms among the queries are compensated, and the resulting update is the unique vector in the sampled subspace whose projections match the observed directional derivatives. As a consequence, each additional query genuinely enlarges and refines the subspace in which the method can make informed progress, rather than merely re-averaging noisy measurements along similar directions. This difference in how query information is used explains why ZO-Align benefits from larger blocks in the strongly convex case, where every improvement in the quality of the update translates into a better contraction factor and hence compounds multiplicatively over iterations. In the limit where the block size equals the dimension and the query matrix has full rank, ZO-Align recovers the true gradient and reduces to a classical full-subspace finite-difference method.
@@
5 High-Dimensional Regime
Modern applications such as large-scale neural network fine-tuning often operate in a regime where the ambient dimension $d$ is extremely large, whereas the query batch size $q$ is constrained to be much smaller than $d$. In this setting, when the columns of $U\in\mathbb{R}^{d\times q}$ are i.i.d. samples from $\mathcal{N}(0,I)$, the Gram matrix $U^{\top}U$ concentrates sharply around a scaled identity:
$U^{\top}U\;\approx\;d\,I_{q}\qquad\text{when }d\gg q,$
so that the sampled directions are nearly mutually orthogonal and $(U^{\top}U)^{-1}$ is, to leading order, a scalar multiple of the identity. A rigorous analysis and empirical verification of the approximation are provided in Appendix A.
This observation clarifies how the benefits of ZO-Align evolve in high dimensions. The main advantage of ZO-Align over simple averaging is that it explicitly accounts for correlations among the sampled directions through the factor $(U^{\top}U)^{-1}$. When $d$ is moderate and $q$ is a nontrivial fraction of $d$, the Gram matrix $U^{\top}U$ can deviate substantially from a scaled identity, and this reweighting meaningfully corrects for the non-orthogonality and unequal norms of the sampled directions. In contrast, when $d\gg q$ and $U^{\top}U\approx dI_{q}$, the sampled directions are already almost orthogonal and have comparable lengths, so the correction degenerates to an almost uniform scaling. As a result, the geometric distinction between ZO-Align and ZO-Avg within the sampled subspace becomes much less pronounced in this extreme high-dimensional regime.
6 Experiments
To complement our theoretical analysis, we evaluate ZO-Avg and ZO-Align under fixed query budgets on both classical objectives and a LLM fine-tuning task. Throughout, we plot performance against the total number of queries, so that the horizontal axis directly reflects zeroth-order computational cost.
6.1 Classical Optimization Problems
We first consider four standard objectives at dimension $d=1000$, representing the regimes studied in Section 4: a strongly convex quadratic, convex logistic regression on synthetic data, the non-convex Rosenbrock function, and a stochastic convex logistic regression objective (details in Appendix C). For each problem we run ZO-Avg and ZO-Align with several query sizes $q$ per iteration under two total budgets, K=20, [^0] K=20{,}000 and $K=500$ (the latter precluding a single full-subspace step). Each configuration is averaged over multiple random seeds.
Across all four problems, the empirical behaviour closely matches our theoretical allocation results. For ZO-Avg, the best performance under a fixed budget is consistently obtained with $q=1$, and increasing $q$ monotonically harms query efficiency, in line with our Theorems. For ZO-Align, larger query blocks yield clear gains in the strongly convex and stochastic convex cases, where the convergence factors depend linearly on $q$, while in the deterministic convex and non-convex problems the curves for different $q$ are much closer, reflecting the allocation-insensitive rates. Overall, these experiments confirm the query allocation dichotomy of Section 4.3: ZO-Avg prefers single-query steps, whereas ZO-Align can exploit larger blocks without the collapse in efficiency seen for ZO-Avg.
Figure 5: Finetuning Qwen3-0.6b on SST2 and CB.
6.2 High-Dimensional Language Model Fine-Tuning
We next examine a high-dimensional setting where the parameter dimension $d$ is extremely large while the feasible query batch size $q$ remains modest ($q\ll d$). We fine-tune the Qwen3-0.6B model on two sentence classification benchmarks, SST-2 and CB, using zeroth-order optimization on the cross-entropy loss. The parameters being updated are treated as a single vector in $\mathbb{R}^{d}$; further protocol details are given in Appendix C.
We compare ZO-Avg with an approximate variant of ZO-Align adapted to the high-dimensional regime. For ZO-Align, forming and inverting the full Gram matrix $U^{\top}U$ is prohibitively expensive at this scale. Motivated by the concentration result in Section 5, we approximate the Gram matrix by its diagonal. We fix a total query budget of K=20, [^0] K=20{,}000 function evaluations and vary the per-iteration query count over $q\in\{1,10,100\}$. A constant step size is tuned separately for each (estimator, $q$) pair by grid search.
As shown on Figure 5, on both SST-2 and CB, ZO-Avg continues to favour small query blocks: the best optimization performance per query is achieved for $q=1$, and increasing $q$ consistently slows progress under the fixed budget. The diagonal ZO-Align variant closely tracks ZO-Avg across all choices of $q$, with the two sets of curves almost indistinguishable at the scale of our plots. This behaviour matches the high-dimensional analysis in Section 5: when $d\gg q$, the Gram matrix concentrates near a scaled identity, so the projection step in ZO-Align reduces essentially to a rescaled version of the ZO-Avg update. In realistic large-scale language-model fine-tuning, the query allocation dichotomy for ZO-Avg therefore remains clearly visible, while the additional geometric structure of ZO-Align provides little practical gain beyond what can be achieved by a suitably tuned ZO-Avg scheme.